OpenStudy (anonymous):

Solve the word problem. The height of a triangle with base b and area A is given by the formula h = . What is the height of a triangle with area 26 cm2 and base 8 cm? A. 1.625 cm B. 3.25 cm C. 3.5 cm D. 6.5 cm

OpenStudy (anonymous):

@itsbribro

OpenStudy (anonymous):

@Clongoria2

OpenStudy (anonymous):

@Kit_Kat_Nat_<3

OpenStudy (anonymous):

@pandagirl811

OpenStudy (anonymous):

As much as id love to help, i'm hopeless with triangles. You might want to check out http://www.wolframalpha.com/ though cause thats what i normally go to when i'm stuck.

OpenStudy (anonymous):

\[area~A=\frac{ 1 }{ 2 }\times base \times height\]

OpenStudy (clongoria2):

So the formula you have is h= 1/2bh. You know that A = 26 = 1/2 8 x H. So then you can just solve this to get the height. 1/2 x 8 = 4. 26 / 4 = 6.5. We can check to see if 6.5 = H. A = 1/2bh 8 x 6.5 = 52. 52 / 2 = 26. A = 26

OpenStudy (anonymous):

thanks